OpenStudy (anonymous):

PLEASE HELP!!! I have a math test tomorrow and need help with substitution and elimination. 1) y=5x+3, y=x-9 2) 2x-y=5, -3x+4y=0

OpenStudy (jdoe0001):

\(\large { (1)\ hint \begin{cases} {\color{brown}{ y}}=5x+3\\ {\color{brown}{ y}}=x-9 \\ \quad \\ {\color{brown}{ y}}={\color{brown}{ y}} \end{cases} }\) you could do the same for the (2) one

OpenStudy (jdoe0001):

or you could just solve the 1st equation of the (2) set for "y" only and substitute on the 2nd equation of (2)

OpenStudy (anonymous):

It'd be nice if you laid out the equation so could follow through it, instead of guessing and getting it wrong

OpenStudy (anonymous):

@jdoe0001

OpenStudy (isaiah.feynman):

That guy jdoe0001 is good. Follow his lead @Syli

OpenStudy (zpupster):

elimination: move all variables to one side y=5x+3 subtract 5x y=x-9 subtract x so, y-5x=3 y-x =-9 lets eliminate y bymullt. here by -1 and then add the 2 equations together so, y-5x =3 -y+x = 9 ____________ -4x = 12 divide both by -4 x=12/-4 or - 3 plug in for x and get y y =-3-9 y=-12 do a check -12-5(-3) =3 -12+15=3 3=3 check here is substitution y=5x+3 y=x-9 since y = x-9 plug in for y in eq(1) x-9= 5x+3 add 9 both sides and subtract 5x so -4x= 12 x=-3 we already found y and did the check good luck on the test.

OpenStudy (anonymous):

Thank you very much! :) @zpupster